# Dimensionality Reduction on Complex Vector Spaces for Euclidean Distance with Dynamic Weights

### v3

@@ 3.1 Tail bounds and Orlicz norms of arbitrary Rademacher chaoses
$\small\begin{split}&\textnormal{Pr}\left[\displaystyle\left|\sum_{i=1}^{k}S_{i}\right|>W(S)\left(c_{1}(\gamma)\sqrt{k}\sqrt{t}+c_{2}(\gamma)t^{\nicefrac{{\gamma}}{{2}}}\right)\right]\\ &\quad\leq\textnormal{Pr}\left[\displaystyle\left|\sum_{i=1}^{k}S_{i}\right|>C_{\nicefrac{{2}}{{\gamma}}}\left\lVert S\right\rVert_{\psi(\nicefrac{{2}}{{\gamma}})}\left(\sqrt{k}\sqrt{t}+2^{\gamma-\nicefrac{{1}}{{2}}}t^{\nicefrac{{\gamma}}{{2}}}\right)\right].\end{split}$
@@ 4.2 Results in high probability
$\quad+2\sum_{i=1}^{k}\sum_{j_{1},j_{2}\in I^{\prime}_{2,d}}\!\textnormal{Re}\left(\displaystyle A_{i,j_{1}}A_{i,j_{2}}^{3}\right)x_{j_{1}}x_{j_{2}}w_{j_{2}}^{2}$
$\quad+\sum_{i=1}^{k}\sum_{j_{1},j_{2}\in I^{\prime}_{2,d}}\!\textnormal{Re}\left(\displaystyle A_{i,j_{1}}^{2}A_{i,j_{2}}^{2}\right)x_{j_{1}}^{2}w_{j_{2}}^{2}$
$\quad+\sum_{i=1}^{k}\sum_{j=1}^{d}w_{j}^{2}x_{j}^{2}$
@@
$\small\begin{split}&\textnormal{Pr}\left[\displaystyle|\rho(g(x),w)-\left\lVert x\right\rVert_{w}^{2}|>t\right]\\ &\quad\leq\begin{cases}8\exp\left(-k\frac{1}{\left(c_{1}8\left\lVert w\right\rVert_{4}^{2}\left\lVert x\right\rVert_{2}^{2}\right)^{2}}t^{2}\right)&\text{if }(\dagger)\\ 8\exp\left(-\sqrt{k}\frac{1}{\sqrt{c_{2}8\left\lVert w\right\rVert_{4}^{2}\left\lVert x\right\rVert_{2}^{2}}}\sqrt{t}\right)&\text{otherwise}\end{cases}\end{split}$
@@ 4.4 Improved results for near-uniform vectors
$\small\begin{split}&\textnormal{Pr}\left[\displaystyle|\bar{\rho}(\bar{g}(x),w)-\left\lVert x\right\rVert_{w}^{2}|>t\right]\\ &\quad\leq\max\begin{cases}8\exp\left(-kc_{1}\frac{L^{3}}{\left(\left\lVert w\right\rVert_{4}^{2}\left\lVert x\right\rVert_{2}^{2}\right)^{2}}t^{2}\right)&\\ 8\exp\left(-\sqrt{k}c_{2}\frac{L^{\nicefrac{{3}}{{4}}}}{(\left\lVert w\right\rVert_{4}^{2}\left\lVert x\right\rVert_{2}^{2})^{\nicefrac{{1}}{{2}}}}\sqrt{t}\right)\end{cases}\end{split}$
@@ Acknowledgments
This work was supported in part by MUR PRIN 20174LF3T8 AHeAD project, by MUR PNRR CN00000013 National Center for HPC, Big Data and Quantum Computing, by Uni-Impresa Big-Mobility project, and by Marsden Fund (MFP-UOA2226).

### v4

@@ 3.1 Tail bounds and Orlicz norms of arbitrary Rademacher chaoses
$\small\begin{split}&\textnormal{Pr}\left[\displaystyle\left|\sum_{i=1}^{k}S_{i}\right|>W(S)\left(c_{1}(\gamma)\sqrt{k}\sqrt{t}+c_{2}(\gamma)t^{\nicefrac{{\gamma}}{{2}}}\right)\right]\\ &\qquad\leq\textnormal{Pr}\left[\displaystyle\left|\sum_{i=1}^{k}S_{i}\right|>C_{\nicefrac{{2}}{{\gamma}}}\left\lVert S\right\rVert_{\psi(\nicefrac{{2}}{{\gamma}})}\left(\sqrt{k}\sqrt{t}+2^{\gamma-\nicefrac{{1}}{{2}}}t^{\nicefrac{{\gamma}}{{2}}}\right)\right].\end{split}$
@@ 4.2 Results in high probability
$\qquad+2\sum_{i=1}^{k}\sum_{j_{1},j_{2}\in I^{\prime}_{2,d}}\!\textnormal{Re}\left(\displaystyle A_{i,j_{1}}A_{i,j_{2}}^{3}\right)x_{j_{1}}x_{j_{2}}w_{j_{2}}^{2}$
$\qquad+\sum_{i=1}^{k}\sum_{j_{1},j_{2}\in I^{\prime}_{2,d}}\!\textnormal{Re}\left(\displaystyle A_{i,j_{1}}^{2}A_{i,j_{2}}^{2}\right)x_{j_{1}}^{2}w_{j_{2}}^{2}$
$\qquad+\sum_{i=1}^{k}\sum_{j=1}^{d}w_{j}^{2}x_{j}^{2}$
@@
$\small\begin{split}&\textnormal{Pr}\left[\displaystyle|\rho(g(x),w)-\left\lVert x\right\rVert_{w}^{2}|>t\right]\\ &\qquad\leq\begin{cases}8\exp\left(-k\frac{1}{\left(c_{1}8\left\lVert w\right\rVert_{4}^{2}\left\lVert x\right\rVert_{2}^{2}\right)^{2}}t^{2}\right)&\text{if }(\dagger)\\ 8\exp\left(-\sqrt{k}\frac{1}{\sqrt{c_{2}8\left\lVert w\right\rVert_{4}^{2}\left\lVert x\right\rVert_{2}^{2}}}\sqrt{t}\right)&\text{otherwise}\end{cases}\end{split}$
@@ 4.4 Improved results for near-uniform vectors
$\small\begin{split}&\textnormal{Pr}\left[\displaystyle|\bar{\rho}(\bar{g}(x),w)-\left\lVert x\right\rVert_{w}^{2}|>t\right]\\ &\qquad\leq\max\begin{cases}8\exp\left(-kc_{1}\frac{L^{3}}{\left(\left\lVert w\right\rVert_{4}^{2}\left\lVert x\right\rVert_{2}^{2}\right)^{2}}t^{2}\right)&\\ 8\exp\left(-\sqrt{k}c_{2}\frac{L^{\nicefrac{{3}}{{4}}}}{(\left\lVert w\right\rVert_{4}^{2}\left\lVert x\right\rVert_{2}^{2})^{\nicefrac{{1}}{{2}}}}\sqrt{t}\right)\end{cases}\end{split}$
@@ Acknowledgments
This work was supported in part by MUR PRIN 2022TS4Y3N EXPAND project, by MUR PNRR CN00000013 National Center for HPC, Big Data and Quantum Computing, by Uni-Impresa Big-Mobility project, and by Marsden Fund (MFP-UOA2226).